Solve the system by system:
8x - 4y = 16
y = 2y - 4
This substitutions have gotten me confused, any help would be appreciated. Thanks

Found 2 solutions by checkley71, kennyxx12:
Answer by checkley71(8403)

Y=2Y-4
Y-2Y=4
-Y=4
Y=-4 SOLUTION NOW SUBSTITUTE -4 FOR Y IN THE OTHERE EQUATION
8X-4(-4)=16
8X+16=16
8X=16-16
8X=0
X=0/8
X=0 SOLUTION.

Answer by kennyxx12(3)

first look at the problem closely y=2y-4
since y equals 2y. minus 4 then y has to be 4
since y+y-4=y
y=4
now enter y to the other equation if substituted to the same formula infinite many solutions will be the answer.this is wrong for this question
8x-4y=16
8x-(4*4)=16
8x-16=16
take the minus 16 to the other side to isolate 8x
and negate the signs to it
8x=16+16
8x=32
divide both sides to isolate x
x=4

THE ANSWER IS=(4,4)